# Transmission Conditions for the Non-Overlapping Schwarz Coupling of Full Order and Operator Inference Models

### v1

@@ 4.2 NO-SAM formulation
A slightly modified version of the OpInf minimization problem in (25) must be solved for the case of Schwarz coupling. In this setting, the subdomain boundary conditions can no longer be assumed constant in time, as is often the case for a monolithic problem. To account for this, we modify (25) for subdomain $i$ as
$\min_{\tilde{~K}_{i},\tilde{~H}_{i},\tilde{~B}_{i}}\sum_{p=1}^{j}\Big{\lVert}\ddot{\hat{~u}}_{i}(t_{p})+\tilde{~K}_{i}\,\hat{~u}_{i}(t_{p})-\tilde{~H}_{i}\,~t_{i}(t_{p})-\tilde{~B}_{i}\,~g_{i}(t_{p})\Big{\rVert}_{2}^{2}+\lambda^{2}\Big{(}\lVert\tilde{~K}_{i}\rVert_{F}^{2}+\lVert\tilde{~H}_{i}\rVert_{F}^{2}+\lVert\tilde{~B}_{i}\rVert_{F}^{2}\Big{)}$ (26)
where $t_{i}(t_{p})$ and $g_{i}(t_{p})$ are altered to explicitly reflect their time dependence. For instance, $g_{1}(t_{p})$ denotes the displacement boundary condition on the boundary $\Gamma$ for subdomain $\Omega_{1}$ at time $t_{p}$. Therefore, snapshot matrices containing boundary condition data on the boundary $\Gamma$ for each timestep are required to infer the reduced order boundary traction and displacement operators.
@@
$\min_{\tilde{~K}_{i},\tilde{~H}_{i},\tilde{~B}_{i}}$ $\sum_{p=1}^{j}\Big{\lVert}\ddot{\hat{~u}}_{i}(t_{p})+\left[\tilde{~K}_{i}+\frac{\beta_{ij}}{\alpha_{ij}}\tilde{~S}_{i}\right]\hat{~u}_{i}(t_{p})-\tilde{~H}_{i}\,~t_{i}(t_{p})-\tilde{~B}_{i}\,~g_{i}(t_{p})-\frac{1}{\alpha_{ij}}\tilde{~R}_{i}\,~c_{i}(t_{p})\Big{\rVert}_{2}^{2}$ (27)
$+\lambda^{2}\Big{(}\lVert\tilde{~K}_{i}\rVert_{F}^{2}+\lVert\tilde{~S}_{i}\rVert_{F}^{2}+\lVert\tilde{~H}_{i}\rVert_{F}^{2}+\lVert\tilde{~B}_{i}\rVert_{F}^{2}+\lVert\tilde{~R}_{i}\rVert_{F}^{2}\Big{)}$
where $\tilde{~S}_{i}$ denotes the Robin stiffness contribution, $\tilde{~R}_{i}$ is the reduced Robin boundary operator, and $c_{i}(t_{p})$ represents the Robin boundary condition data prescribed on $\Gamma$ at time $t_{p}$. The constants $\alpha_{ij}$ and $\beta_{ij}$ have been factored out of the operators to remove their dependence on these parameters. For all relevant numerical examples, the Robin boundary condition data $c_{i}(t_{p})$ is scaled by a factor of $\frac{1}{\sigma_{\max}}$ to enhance stability.
We emphasize that in our OpInf NO-SAM algorithm, the reduced operators appearing in (27) are learned once in the offline stage of the model reduction procedure, similar to the computation of the POD basis ${\bf\Phi}_{r}$. Once these operators have been solved for, we can apply the online stage of the algorithm, in which time-stepping and subdomain synchronization via the NO-SAM algorithm is performed. Specifically, in each time interval $I_{n}=[t_{n},t_{n+1}]$, every Schwarz iteration involving the OpInf model is carried out according to the following steps:
@@
    Project the previous full order displacement, velocity, and acceleration snapshots into the reduced order basis using $\Phi_{r}$.
@@
    Check convergence using (10). If satisfied, advance to the next timestep ($n\leftarrow n+1$); otherwise, move to the next subdomain and return to step 1.
@@ 4.3 Numerical results
Table 2 reports the error as defined in (15), the mean number of Schwarz iterations required for convergence, and the CPU time for each FOM-OpInf model combination using both alternating Dirichlet-Neumann and Robin-Robin transmission conditions. The FOM-FOM results from Table 1 for both transmission condition types are included as a reference. For cases employing a subdomain-local OpInf model, the number of modes used to construct the corresponding POD basis is listed. Here, $M_{1}$ denotes the number of POD modes used for the left domain, and $M_{2}$ the number of POD modes used for the right domain. For POD bases retaining $99.9\%$ of the snapshot energy, it is clear that the inclusion of a subdomain OpInf model consistently increases both the number of Schwarz iterations required for convergence as well as the resulting error. To assess whether this behavior is due to insufficient modal content, we also consider a case with 34 and 29 POD modes for the left and right subdomains, respectively, corresponding to $99.999999\%$ energy retention in each subdomain.
@@
Fig. 6: OpInf-FOM coupling: plots of the NO-SAM displacement, velocity, and acceleration solutions obtained using alternating Dirichlet-Neumann transmission conditions, compared to the monolithic solution (dashed line). Here, the blue subdomain employs an OpInf model and uses a Dirichlet transmission condition, and the red subdomain employs a full order model and uses a Neumann transmission condition.
Similar trends are observed for the Robin–Robin case, where the large errors present in the lower modal content setting generally diminish as the modal content is increased. One exception is observed in the FOM-OpInf case, where the simulation becomes unstable. This instability is likely due to insufficient regularization, as the optimal regularization parameter $\delta$ may vary with the problem setting and model configuration. Among all tested cases, the OpInf-OpInf configuration with Robin–Robin transmission conditions and higher modal content yielded the best overall performance in terms of accuracy, iteration count, and CPU time. This simulation, illustrated in Figure 7, converged in two Schwarz iterations for every timestep, achieving a speedup of $1.77\times$ and an accuracy improvement of $2.11\times$ compared to the reference FOM-FOM result. Recall that the $\alpha_{ij}$ and $\beta_{ij}$ parameters used for the Robin–Robin transmission condition were taken directly from the lowest-error parameter set identified in the FOM-FOM parameter sweep. It would, however, be interesting to generate a similar Pareto plot as in Figure 2 when coupling full order and reduced order models, to determine whether a similar trend emerges or if an even more optimal parameter set exists.
@@
Fig. 7: OpInf-OpInf coupling: plots of the NO-SAM displacement, velocity, and acceleration solutions obtained using Robin-Robin transmission conditions (lowest error parameter set, $\left[\overline{\alpha}_{12},\overline{\alpha}_{21},\beta_{12},\beta_{21}\right]=\left[10^{-3},10^{-3},1,1\right]$), compared to the monolithic solution (dashed line). Here, both the blue and red subdomains employ an OpInf model.
The final numerical example examines the accuracy and stability of the OpInf model within the predictive regime. The problem setting defined in Section 2.1 is used, with the simulation extended to $t_{f}=10^{-2}\,s$, corresponding to five periods of wave propagation along the bar. The corresponding monolithic displacement solution is used to collect 4000 snapshots over the original time interval, which spans half a period of wave propagation. The four configurations considered correspond to the combinations of alternating Dirichlet-Neumann and Robin-Robin transmission conditions with either full order or OpInf reduced order models.
@@ 5 Conclusions and future work
Looking ahead, this study lays the groundwork for extending NO-SAM to predictive regimes, multi-dimensional problems, and large-scale computational frameworks. In particular, integration into three-dimensional (3D) finite element codes such as Norma.jl [26] and deployment within production engineering codes such as SIERRA/SM [2] will provide valuable opportunities for assessing scalability and practical impact. Moreover, the development of optimized domain decomposition strategies and hybrid FOM–ROM model libraries will be essential to fully harness the flexibility and efficiency of NO-SAM. Finally, we plan to explore in future work the SAM-based coupling of alternate non-intrusive ROMs, e.g., ROMs constructed using reproducing kernel Hilbert spaces [9, 10]. Together, these directions position the method as a promising tool for advancing scalable, non-intrusive coupling in computational mechanics and beyond.

### v2

@@ 4.2 NO-SAM formulation
A modified version of the OpInf minimization problem in (24) must be solved in the context of Schwarz coupling. In this case, each subdomain is associated with its own set of operators. Accordingly, the formulation in (24) is adapted for each subdomain $k$ as
$\min_{\tilde{\mathord{\btensor K}}_{k},\tilde{\mathord{\btensor H}}_{k},\tilde{\mathord{\btensor B}}_{k}}\sum_{p=1}^{P}\Big\lVert\ddot{\hat{\mathord{\btensor u}}}_{k}(t_{p})+\tilde{\mathord{\btensor K}}_{k}\,\hat{\mathord{\btensor u}}_{k}(t_{p})-\tilde{\mathord{\btensor H}}_{k}\,\mathord{\btensor t}_{k}(t_{p})-\tilde{\mathord{\btensor B}}_{k}\,\mathord{\btensor g}_{k}(t_{p})\Big\rVert_{2}^{2}+\lambda^{2}\Big(\lVert\tilde{\mathord{\btensor K}}_{k}\rVert_{F}^{2}+\lVert\tilde{\mathord{\btensor H}}_{k}\rVert_{F}^{2}+\lVert\tilde{\mathord{\btensor B}}_{k}\rVert_{F}^{2}\Big)$ (25)
Where, for example, $\hat{\mathord{\btensor u}}_{k}(t_{p})$ denotes the reduced displacement within subdomain $\Omega_{k}$ at time $t_{p}$. For the case of Schwarz coupling, additional snapshot data containing the transmission conditions $\mathord{\btensor g}$ and $\mathord{\btensor t}$ on the interface $\Gamma$ at each time step is required. Only the operators required by the selected transmission condition for each subdomain are learned.
@@
$\min_{\begin{subarray}{c}\tilde{\mathord{\btensor K}}_{k},\tilde{\mathord{\btensor S}}_{k},\tilde{\mathord{\btensor H}}_{k},\\ \tilde{\mathord{\btensor B}}_{k},\tilde{\mathord{\btensor R}}_{k}\end{subarray}}$ $\sum_{p=1}^{P}\Big\lVert\ddot{\hat{\mathord{\btensor u}}}_{k}(t_{p})+\left[\tilde{\mathord{\btensor K}}_{k}+\tfrac{\beta_{ij}}{\alpha_{ij}}\tilde{\mathord{\btensor S}}_{k}\right]\hat{\mathord{\btensor u}}_{k}(t_{p})-\tilde{\mathord{\btensor H}}_{k}\,\mathord{\btensor t}_{k}(t_{p})-\tilde{\mathord{\btensor B}}_{k}\,\mathord{\btensor g}_{k}(t_{p})-\tfrac{1}{\alpha_{ij}}\tilde{\mathord{\btensor R}}_{k}\,\mathord{\btensor c}_{k}(t_{p})\Big\rVert_{2}^{2}$ (26)
$+\lambda^{2}\Big(\lVert\tilde{\mathord{\btensor K}}_{k}\rVert_{F}^{2}+\lVert\tilde{\mathord{\btensor S}}_{k}\rVert_{F}^{2}+\lVert\tilde{\mathord{\btensor H}}_{k}\rVert_{F}^{2}+\lVert\tilde{\mathord{\btensor B}}_{k}\rVert_{F}^{2}+\lVert\tilde{\mathord{\btensor R}}_{k}\rVert_{F}^{2}\Big)$
where $\tilde{\mathord{\btensor S}}_{k}$ denotes the reduced Robin stiffness operator, $\tilde{\mathord{\btensor R}}_{k}$ is the reduced Robin boundary operator, and $\mathord{\btensor c}_{k}(t_{p}):=\alpha_{ij}\,\mathord{\btensor t}_{k}(t_{p})+\beta_{ij}\,\mathord{\btensor g}_{k}(t_{p})$ represents the Robin transmission condition data prescribed on $\Gamma$ for subdomain $k$ at time $t_{p}$. The constants $\alpha_{ij}$ and $\beta_{ij}$ have been factored out of each operator, thereby removing their dependence on these parameters. By learning $\tilde{\mathord{\btensor S}}_{k}$ and $\tilde{\mathord{\btensor K}}_{k}$ independently, one can conveniently tune $\alpha_{ij}$ and $\beta_{ij}$ when necessary. For all relevant numerical examples, the Robin boundary condition data $\mathord{\btensor c}_{k}(t_{p})$ is scaled by a factor of $\frac{1}{\sigma_{\max}}$ to enhance stability.
We emphasize that in our OpInf NO-SAM algorithm, the reduced operators appearing in (26) are learned once in the offline stage of the model reduction procedure, similar to the computation of the POD basis ${\bf\Phi}_{r}$. Once these operators have been solved for, we can apply the online stage of the algorithm, in which time-stepping and subdomain synchronization via the NO-SAM algorithm is performed. Specifically, in each time interval $I_{n}:=[t_{n},t_{n+1}]$, every Schwarz iteration involving the OpInf model is carried out according to the following steps:
@@
    Project the previous full order displacement, velocity, and acceleration snapshots into the reduced order basis using r.
@@
    Check for convergence using (10). If satisfied, advance to the next time step ($n\leftarrow n+1$); otherwise, move to the next subdomain and return to step 1.
@@ 4.3 Numerical results
Table 2 reports the error as defined in (15), the mean number of Schwarz iterations per time step required for convergence, and the CPU time for each FOM-OpInf model combination using both alternating Dirichlet-Neumann and Robin-Robin transmission conditions. The FOM-FOM results from Table 1 for both transmission condition types are included as a reference. For cases employing a subdomain-local OpInf model, the number of modes used to construct the corresponding POD basis is listed. Here, $M_{1}$ denotes the number of POD modes used for the left domain, and $M_{2}$ the number of POD modes used for the right domain. For POD bases retaining $99.9\%$ of the snapshot energy, it is clear that the inclusion of a subdomain OpInf model consistently increases both the number of Schwarz iterations required for convergence as well as the resulting error. To assess whether this behavior is due to insufficient modal content, we also consider a case with 34 and 29 POD modes for the left and right subdomains, respectively, corresponding to $99.999999\%$ energy retention in each subdomain.
@@
Fig. 6: OpInf-FOM coupling: plots of the NO-SAM displacement, velocity, and acceleration solutions obtained using alternating Dirichlet-Neumann transmission conditions, compared to the monolithic solution (dashed line). Here, the blue subdomain $\Omega_{1}$ employs an OpInf model and uses a Dirichlet transmission condition, and the red subdomain $\Omega_{2}$ employs a full order model and uses a Neumann transmission condition.
Similar trends are observed for the Robin–Robin case, where the large errors present in the lower modal content setting generally diminish as the modal content is increased. One exception is observed in the FOM-OpInf case, where the simulation becomes unstable. This instability is likely due to insufficient regularization, as the optimal regularization parameter $\delta$ may vary with the problem setting and model configuration. Among all tested cases, the OpInf-OpInf configuration with Robin–Robin transmission conditions and higher modal content yielded the best overall performance in terms of accuracy, iteration count, and CPU time. This simulation, illustrated in Figure 7, converged in two Schwarz iterations at every time step, achieving a speedup of $1.77\times$ and an accuracy improvement of $2.11\times$ compared to the reference FOM-FOM result. Recall that the $\alpha_{ij}$ and $\beta_{ij}$ parameters used for the Robin–Robin transmission condition were taken directly from the lowest-error parameter set identified in the FOM-FOM parameter sweep. It would, however, be interesting to generate a similar Pareto plot as in Figure 2 when coupling full order and reduced order models, to determine whether a similar trend emerges or if an even more optimal parameter set exists.
@@
Fig. 7: OpInf-OpInf coupling: plots of the NO-SAM displacement, velocity, and acceleration solutions obtained using Robin-Robin transmission conditions (lowest error parameter set, $\left[\overline{\alpha}_{12},\overline{\alpha}_{21},\beta_{12},\beta_{21}\right]=\left[10^{-3},10^{-3},1,1\right]$), compared to the monolithic solution (dashed line). Here, both the blue and red subdomains $\Omega_{1}$ and $\Omega_{2}$ employ an OpInf model.
The final numerical example examines the accuracy and stability of the OpInf model within the predictive regime. The problem setting defined in Section 2.1 is used, with the simulation extended to $t_{f}=10^{-2}\,s$, corresponding to five periods of wave propagation along the bar. The corresponding monolithic displacement solution is used to collect 4000 snapshots over the original time interval, which spans half a period of wave propagation. The four configurations considered correspond to the combinations of alternating Dirichlet-Neumann and Robin-Robin transmission conditions with either full order or OpInf reduced order models. Here, both OpInf-OpInf configurations correspond to the higher modal content setting, with $M_{1}=34$ and $M_{2}=29$.
@@ 5 Conclusions and future work
Looking ahead, this study lays the groundwork for extending NO-SAM to more complex predictive regimes, multi-dimensional problems, and large-scale computational frameworks. In particular, integration into three-dimensional (3D) finite element codes such as Norma.jl [25] and deployment within production engineering codes such as SIERRA/SM [2] will provide valuable opportunities for assessing scalability and practical impact. Moreover, the development of optimized domain decomposition strategies and hybrid FOM–ROM model libraries will be essential to fully harness the flexibility and efficiency of NO-SAM. Finally, we plan to explore in future work the SAM-based coupling of alternate non-intrusive ROMs, e.g., ROMs constructed using reproducing kernel Hilbert spaces [9, 10]. Together, these directions position the method as a promising tool for advancing scalable, non-intrusive coupling in computational mechanics and beyond.
@@ Acknowledgments
Appendix
Here, we derive the additional stiffness and boundary force operators arising from the Robin transmission conditions, as expressed in (11). Without loss of generality, consider the subdomain problem in $\Omega_{1}$. We begin with the force vector from the subdomain-local semi-discrete formulation, as given in the first line of (6). Here, we only consider the force contribution due to the transmission condition at the interface, denoted $\mathord{\btensor f}^{(s+1)}_{1}$ and defined as
$\mathord{\btensor f}^{(s+1)}_{1}:=\int_{\Gamma}\mathord{\btensor T}^{(s+1)}_{1}\cdot\boldsymbol{\xi}\ dS$ (27)
It is important to note that this term is only integrated over the subdomain interface $\Gamma$. Solving for $\mathord{\btensor T}^{(s+1)}_{1}$ in the third line of (6) and substituting the result into (27), while using the expression for $^{(s+1)}_{1}$ and assuming no relaxation ($\theta_{1}=1$), yields
$\mathord{\btensor f}^{(s+1)}_{1}=\int_{\Gamma}\left[\frac{1}{\alpha_{12}}\mathord{\btensor c}_{1}^{(s)}-\frac{\beta_{12}}{\alpha_{12}}\mathord{\btensor u}_{1}^{(s+1)}\right]\cdot\boldsymbol{\xi}\ dS$ (28)
where $\mathord{\btensor c}_{1}^{(s)}:=\alpha_{12}\mathord{\btensor T}_{2}^{(s)}+\beta_{12}\mathord{\btensor u}_{2}^{(s)}$ denotes the Robin transmission condition prescribed by the neighboring subdomain. The force term in (28) can then be decomposed into two separate integrals:
$\mathord{\btensor S}^{(s+1)}_{1}:=\frac{\beta_{12}}{\alpha_{12}}\int_{\Gamma}\mathord{\btensor u}_{1}^{(s+1)}\cdot\boldsymbol{\xi}\,dS$ (29)
$\mathord{\btensor R}^{(s+1)}_{1}:=\frac{1}{\alpha_{12}}\int_{\Gamma}\mathord{\btensor c}_{1}^{(s)}\cdot\boldsymbol{\xi}\,dS$ (30)
Together, these expressions, when substituted into the external force vector of the first line of (6), yield the final semi-discrete matrix-vector form. Here, $\mathord{\btensor S}^{(s+1)}_{1}$ contributes an additional stiffness component, while $\mathord{\btensor R}^{(s+1)}_{1}$ contributes an additional external force component. The remaining $\mathord{\btensor F}^{(s+1)}_{1}$ vector in (11) represents the external body force together with any applied traction on the outer domain boundary $\partial_{T}\Omega$. Observe that the negative sign in (28) disappears when the second term is moved to the left-hand side of (11).